### v1

# Citation: Kuric E., Demcak P., Krajcovic M. (2025) Talking Surveys: How Photorealistic Embodied Conversational Agents Shape Response Quality, Engagement, and Satisfaction

Abstract
To keep card sorting with a lot of cards concise, a common strategy for gauging mental models involves presenting participants with fewer randomly selected cards instead of the full set. This is a decades-old practice, but its effects lacked systematic examination. To assess how randomized subsets affect data, we conducted an experiment with 160 participants. We compared results between full and randomized 60% card sets, then analyzed sample size requirements and the impacts of individual personality and cognitive factors. Our results demonstrate that randomized subsets can yield comparable similarity matrices to standard card sorting, but thematic patterns in categories can differ. Increased data variability also warrants larger sample sizes (25-35 for 60% card subset). Results indicate that personality traits and cognitive reflection interact with card sorting. Our research suggests evidence-based practices for conducting card sorting while exposing the influence of study design and individual differences on measurement of mental models.
Keywords: card sorting, mental models, personality traits, cognitive reflection, randomized subsets
1 Introduction
Nowadays, online card sorting tools represent a convenient and modern medium for administering card sorting [Melissourgos and Katsanos, 2021, Paladino et al., 2017, Righi et al., 2013, Bussolon et al., 2006]. Card sorting is a classic, structured method of user research, used to gain understanding of mental models, i.e., how users mentally organize and label various concepts or content [Jansen et al., 2023, Doubleday, 2013, Tchivi et al., 2025, Jean et al., 2018, Lewis and Hepburn, 2010]. Online tools are popular because they can facilitate the method remotely and without a moderator, as well as provide advanced analyses, metrics and visualizations for easier data interpretation [Righi et al., 2013, Wentzel et al., 2016, Thomas and Johnson, 2013]. The primary purpose of card sorting in user research is to inform the design of information architecture—systems for organizing content and navigation structures [Tchivi et al., 2025, Ding et al., 2017, Schall, 2014].
In spite of its prevalence however, online card sorting faces challenges that stem from gaps in methodological understanding. The duration and complexity of card sorting can increase significantly when participants create categories from scratch (i.e., open card sorting) and the number of cards to sort is high. This contributes to fatigue and decreases participant engagement, a critical risk for human-centered research as it discourages interaction [Jones et al., 2025, Doherty and Doherty, 2018]. Decreased interest can also be linked to satisficing [Blazek and Siegel, 2024], manifested as groupings that are viewed as good-enough rather than the best [Hudson, 2007].
A common strategy implemented in online card sorting tools (e.g., UXtweak [^1] User research platform UXtweak: https://www.uxtweak.com/) to reduce the cognitive burden of too many cards is to present each participant with fewer randomly selected cards from a larger set. However, research that would rigorously and empirically investigate the validity of this approach is limited. Aside from reducing statistical significance (individual cards are seen by fewer participants as part of different combinations), different selections could introduce biases into data by affecting participants’ sorting criteria [Rugg and McGeorge, 2005]. Most claims suggesting the approach [Katsanos et al., 2023, Wood and Wood, 2008, Tchivi et al., 2025] originate from a sole poster presented at the Annual Meeting of the Usability Professionals Association—which, to the best of our knowledge, is not publicly available, nor was it subject to methodological peer review [Tullis and Wood, 2005].
To address this issue, we conduct a comprehensive study with the primary goal of comparing the results of card sorting depending on whether participants sorted a full list of cards or its subset. As a sub-issue, randomized subsets in card sorting may have negative impacts on the sample size needed to obtain reliable results. To obtain empirical findings about optimal sample size, we explore the relationship between the number of participants and sorting results. As a second sub-issue, the natural diversity in participants’ psychological profiles—personality or cognitive reflection—may introduce variability detrimental for the use of randomized subsets. Representation of specific profiles may also be beneficial, leading us to investigate the effects of personality and cognitive reflection. This is a topic yet to be examined for card sorting in general, in spite of personality and cognitive reflection having established effects on user behavior, experience and perceived usability [Hattie Wilczewski and Kortum, 2024, Liapis et al., 2019, Özbek et al., 2014, Mosleh et al., 2021], as well as relationships with mental models in other contexts [Vˆırgă et al., 2014, Fisher et al., 2012].
Therefore, key contributions of this research consist of:
  • design, administration and data analysis of an experiment with 160 participants, involving card sorting with 2 card sets by 50 cards (presented either fully or as a 30-card subset) and assessment of participants’ personality and cognitive reflection,
  • validation of card sorting with randomized card subsets, discovering similar card pairing patterns, but different distribution of category themes from standard card sorting,
  • sample size determination formula that accounts for size of the randomized subset to improve the consistency of card sorting results,
  • explorative analysis demonstrating diverse effects of Big Five personality traits and Cognitive Reflection on both standard and randomized subset card sorting.
This article has the following structure. Section 2 elaborates further upon concepts relevant to understanding the background and previous works. The experimental method as well as the design and implementation of the card sorting tool is described in Section 3. Our findings are analyzed in Section 4, then interpreted and discussed in Section 5, which also covers our study’s limitations and proposes future research. Section 6 serves as the conclusion.
2 Background and Related Work
Among established card sorting techniques, open card sorting is one of the most frequently used [Katsanos et al., 2023]. As a generative activity, it allows for a high degree of expression and variety since participants are free to create categories by their own logic [Sione Paea and Bulivou, 2024]. Closed or hybrid card sorting evaluate pre-existing categories, while tree testing is a complementary method for testing information architecture hierarchies [Kuric et al., 2025] designed based on card sorting. We focus on open card sorting because of its prevalence and analytical richness. The optimal sample size for card sorting was determined by Lantz et al. [2019] as between 10 and 15 participants, a notion we seek to explore further in the context of sorting with randomized subsets.
In this section, we further illustrate potential links of personality and cognitive reflection to card sorting, and their relevance to the problem of open card sorting with randomized subsets
2.1 Big Five personality model
At present, personality—the persistent characteristics of a person, which can manifest through behavior, motivations, thoughts and preferences [Fajkowska, 2018]—can be represented by various models. The Big Five model emerged as a unifying taxonomy that is still well-suited for thorough personality measurement [Feher and Vernon, 2021, Anglim and O’connor, 2019]. Even though alternatives exist, such as the six-factor HEXACO—which offers higher orthogonality along with potential theoretical and practical advantages [Ashton and Lee, 2007, Thielmann et al., 2022]—the Big Five remains the most prolific personality model at use today.
Big Five consists of five traits: Negative Emotionality, Extraversion, Agreeableness, Conscientiousness and Open-Mindedness, associated with the following characteristics [Zekry and McKee, 2023, Tatnall, 2020]:
  • Negative Emotionality. Susceptibility or resistance to negative feelings like fear or stress.
  • Extraversion. Sociability, liveliness, preference for group company.
  • Agreableness. Altruism, trust, cooperation, empathy.
  • Conscientiousness. Sense of duty, organization, ambition.
  • Open-mindedness. Acceptance to new concepts and experiences, imagination.
The relevance of these traits for card sorting can be judged by their implications. For example, high Open-Mindedness could reasonably make categories more creative, while low Agreeableness could result in superficial categories due to uncooperativeness and low cognitive effort. Insights about the potential effects of these traits could have applications for research personalization. Recruitment sources may also be biased toward certain personalities [Rife et al., 2016], such as crowdsourcing platforms with higher representation of introverts than the wider population [Burnham et al., 2018]. Therefore, if the influence of personality is significant, it could make achieving generalizable results more challenging [Mullen et al., 2021].
2.2 Cognitive Reflection
Cognitive reflection is a thinking disposition, introduced by Frederick [2005] as a tendency toward reflective (System 2) rather than intuitive (System 1) thinking. It measures the ability to find correct answers through analytical reasoning rather than being swayed by intuitive heuristics [Shtulman and Young, 2023]. It also has a broader significance in decision making, reflecting the subject’s ability to apply better heuristics in appropriate contexts [Campitelli and Labollita, 2010, Cokely and Kelley, 2009].
Cognitive reflection is linked to constructs of cognitive abilities such as intelligence and numeracy or executive functions [Blacksmith et al., 2019, Liberali et al., 2012]. However, it was found by research to be a better predictor of performance in various types of mental tasks [Toplak et al., 2011, Campitelli and Gerrans, 2014]. The capacity for active open-minded thinking can augment intuition, although there is a limit determined by task complexity [Thompson et al., 2018]. In card sorting tasks where no answers are objectively correct, high cognitive reflection might result in intuitively more situation- and environment-aware mental models. Suppression of intuition might also result in mental models based more strongly on reflective reasoning.
3 Study Method
For our assessment of using randomized card subsets in card sorting, we designed our experiment as a standard comparative study where card sorting is administered under two conditions: with the full list of cards as is the norm, and a subset of cards as the alternative. To prevent carryover effects between conditions, the experiment used between-subject design, thus requiring a sample of larger size. The inherent tradeoff of this approach—threat to internal validity due to comparing different groups—was mitigated through random stratified sampling, as shown in 3.1 Participants. To increase overall external validity, our method involved parallel evaluation of card sorting variants with two card sets from two distinct domains. These card sets (see 3.3 Materials for further details) represent different levels of conceptual complexity and familiarity to participants.
The research questions were derived directly from the primary goal of our study, as well as its sub-issues (see 1 Introduction for their justification):
  • RQ1: How does card sorting with a randomized subset of cards affect its results?
  • RQ2: How does the number of participants in card sorting with a randomized subset of cards affect its similarity to sorting with the full card set?
  • RQ3: How do participants’ personality traits affect their card sorting results?
  • RQ4: How does participants’ cognitive reflection affect their card sorting results?
3.1 Participants
In total, 160 participants completed the experiment, 40 in each of four experimental conditions. Recruitment implemented a random stratified sampling strategy, performed through UXtweak’s User Panel [^2] UXtweak’s User Panel: https://www.uxtweak.com/user-participant-recruitment to make use of its large size and pre-screening attributes. All participants were residents of the United Kingdom. The overall sample, as well as the four individual sub-groups were balanced in terms of gender, $\chi^{2}(3,160)=0.08,p=.99$. Their age distributions ($\chi^{2}(9,160)=0.17,p=1.0$) were set to approximately match the population of internet users [^3] Distribution of internet users: https://www.statista.com/statistics/272365/age-distribution-of-internet-users-worldwide/. Other characteristics, such as education, $\chi^{2}(6,160)=5.94,p=.4$ and income $\chi^{2}(12,160)=7.32,p=.84$ also had similar distributions across variants. Post-hoc tests signal that participants’ frequency of visiting websites from the domain relevant to the card sorting was consistent within groups: e-commerce $\chi^{2}(2,80)=0.29,p=.87$, and banking $\chi^{2}(2,80)=3.25,p=.20$.
To determine the sample size a priori, G*Power was used. The analysis was based on a two-tailed Wilcoxon-Mann-Whitney test with 80% power, an alpha level of .05, and a Cohen’s $d$ of 0.65 (approximately equivalent to a medium effect size of $r=0.3$ to capture meaningful actionable effects). This indicated that a total of 80 participants would be sufficient. Because the two testing conditions (e-commerce and banking, see 3.2 Procedure) were to be analyzed separately, a total of 160 participants were acquired.
3.2 Procedure
Participants completed a study in the Card Sorting tool which was thus aligned with the tools standard procedural framework (see 3.3.1 Card Sorting tool). The pre-study questionnaire contained a demographic questionnaire to monitor the sample’s distribution and a Cognitive Reflection Test (CRT) to measure cognitive capacity. The post-study questionnaire contained the personality assessment questionnaire (BFI-2-XS) [Soto and John, 2017a]. Both questionnaires also incorporated an attention check question to verify participants’ engagement and effort during the study. This helped us identify participants who may exhibit straight-lining behavior in the online environment [Krajcovic et al., 2025].
The primary card sorting activity employed a different set of cards depending on one of four experimental conditions: Full-E, Subset-E, Full-B and Subset-B as shown in Table 1. There were two card sets from two domains—ecommerce and banking—presented either as the full set, or its randomized subset.
Table 1: Experimental variants utilized in the study differing in the number of cards and utilized domain.
Variant Card set (size) Domain
Full-E All (50) e-commerce
Subset-E Random subset (30) e-commerce
Full-B All (50) banking
Subset-B Random subset (30) banking
3.3 Materials
Collection of card sorting data demanded a tool for online administration, alongside the design card sets. Psychometric methods were incorporated to measure the mental constructs investigated for their potential effects on card sorting: Big Five personality traits and Cognitive Reflection.
3.3.1 Card Sorting tool
For conducting card sorting research through a web browser, we used Card Sorting, a tool from the online user experience research platform UXtweak [^4] User research platform UXtweak: https://www.uxtweak.com/. Card Sorting provides a flexible environment for investigating design factors of card sorting studies (e.g., card subset randomization, sample size), collection of data (card sorting results, personality assessment questionnaire, cognitive reflection test) and continuous improvement based on study implications.
From the participant’s perspective, a Card Sorting study is opened through a link. The exact design of a study is contingent on its configuration by its conductor. Individual steps are customizable or optional, but follow a consistent order. For ethical and legal reasons, informed consent precedes the collection of data. Initial steps include a welcome message, screening and instructions. A pre-study questionnaire can be used with a variety of open-text and close-ended questions (e.g., multiple choice, Likert scale). Additionally, a questionnaire can also be placed after the main card sorting activity. There, it can function as debriefing, or postpone questions that could prime participants, such as by giving away context that could skew the results.
The core card sorting activity is designed to maximize engagement, promoting dynamic interactions where cards can be moved around and categories can be created and changed intuitively and quickly. This is achieved through a skeuomorphic interaction design that resembles original paper-based card sorting. Users can drag and drop cards from the panel on the left to the space on the right, which contains categories (see Figure 1). New categories can be created by dragging cards over empty space. They can be named by clicking their header.
Figure 1: Card Sorting user interface during the Card Sorting activity. Unsorted cards are on the left while categories containing cards are on the right. Example demonstrates an in-progress sorting of electrical devices.
During study configuration, researchers can create or import any number of cards. To lower the cognitive load on participants, there is an option to present each individual with a random subset of fixed size. To obtain an appropriately sized sample, participants can be recruited and screened either from external channels, or internally through UXtweak’s User Panel service. The panel provides targeting of specific audiences based on demographic, technical and other attributes.
Additionally, the researcher-facing part of the Card Sorting tool provides comprehensive analytical capabilities that support the exploration of user’s mental models through measures, data analysis and visualizations. Information and methods relevant for comparison of experimental variables in card sorting are discussed in 3.4 Measures and 3.5 Data analysis.
3.3.2 Card sets
Determining the number of cards in the card sets was key to our experimental design. Due to lack of universally accepted standards and empirical validation, most practical recommendations about the number of cards in card sorting are based on anecdotal evidence [Tchivi et al., 2025, Spencer, 2009]. The full list of cards was to be sorted in two experimental conditions as a source of ground truth. Therefore, 50 was chosen as a reasonable number to help preserve engagement [Spencer, 2009]. With 50 cards in total, 30 cards also represent the subset size, which doubles as 60% of the total as suggested by literature [Tchivi et al., 2025] and the bottom recommended limit to ensure the emergence of patterns [Spencer, 2009].
Two card sets were designed to represent different websites where card sorting can be conducted in practice—e-commerce and banking. Cards in the e-commerce variant contain electrical devices and appliances—physical objects that most humans should recognize and be able to sort. All products are of a specific type to maintain a reasonable degree of challenge as opposed to sorting products into broad categories like fashion, electronics, food, etc. The cards in the banking card set, representing intangible concepts, are more abstract. They are also less familiar to the average person, requiring a small degree of financial literacy. As such, the banking card sort is more difficult.
Card labels for both sets were inspired by existing websites. We reviewed information architectures of 5 existing English e-commerce/banking websites, identifying common labels and patterns, as well as cards that could be realistically challenging to sort. Thematically, cards were designed to fit into approximately 5-8 categories. The complete sets are available in the appendix (see Appendix A).
3.3.3 Big Five Inventory-2
To minimize potential effects of questionnaires on engagement and thus on user behavior during card sorting, we used the extra short form of the Big Five Inventory-2 (BFI-2-XS) [Soto and John, 2017a] to assess Big Five personality traits. While the standard BFI-2 is a robust instrument for measuring the five personality traits and their respective three facets [Soto and John, 2017b], it contains 60 Likert scale items, which is considerable when the efficiency of data collection is a concern. Card sorting itself is already an effort-intensive process. As such, the 15-item BFI-2-XS was chosen to prevent fatigue from participants. In spite of its short length, BFI-2-XS captures the Big Five traits—Extraversion, Conscientiousness, Open-Mindedness, Agreeableness and Negative Emotionality—with 80% of the variability of BFI-2 [Soto and John, 2017a].
3.3.4 Cognitive Reflection Test
Cognitive reflection was measured through the standard Cognitive Reflection Test (CRT) [Brañas-Garza et al., 2019, Frederick, 2005]. This questionnaire comprises 3 questions, realized in close-ended form to facilitate their analysis [Sirota and Juanchich, 2018]. The questionnaire’s capacity to predict cognitive reflection is resistant even to repeated administrations, which Bialek and Pennycook [2018] attributed to only cognitively reflective individuals having the ability to independently reflect on their error. The options were ordered randomly to ensure unpredictable placements of the right answers.
3.3.5 Feedback questionnaire
In addition to comparing objective characteristics of card sorting results, we also aimed to assess participants’ attitudes and perceptions. As such, we designed a 5-point Likert-scale questionnaire for measuring key aspects of subjective card sorting experience, as shown in Table 2. A single open-ended question for general feedback was also included.
Table 2: Feedback questionnaire for measuring perception of card sorting.
Perceived construct Question From (1) To (5)
Label clarity How clear or unclear was it to understand the card labels? Very unclear Very clear
Difficulty How easy or difficult was it to categorise the cards? Very difficult Very easy
Focus How did you find it to maintain concentration while sorting cards? Very challenging Very easy
Time spent How would you rate the amount of time it took you to complete the card sorting activity? Much too long Much too short
Number of cards What’s your opinion of the number of cards that you were asked to sort? Too many cards Too few cards
3.4 Measures
Since our research questions explore the impacts of a multitude of factors on card sorting results, our measures represent typical constructs involved in card sorting analysis. Open card sorting results assessed for their similarity comprise user-generated categories, their labels, and placements of cards assigned categories.
Behavioral indicators captured during card sorting include standard measures such as the time spent on the card sorting activity and the number of categories created [Righi et al., 2013]. These measures could be expected to increase automatically with the increased number of cards, however the specifics of the relationship are open for exploration (e.g., in spite of more cards, the number of categories may not increase if participants are satisfied after creating a specific model, some personalities may be more inclined to split larger categories). The attitudinal perspective includes perceived variables: Time Spent, Card Count, Label Clarity, Difficulty and Focus, determined as ratings from the post-study questionnaire Likert scales.
3.5 Data analysis
The Shapiro–Wilk test was performed for variables including time spent, number of created categories, Big Five scores or CRT scores, discovering that they do not conform to a normal distribution. Therefore, the nonparametric Mann–Whitney U test was used as the primary statistical method. For in-depth comparative analysis of results, we delved into card sorting data from three perspectives: structural, linguistic, and thematic.
Chi-squared tests were performed for categorical variables. To assess the effects of sample size (RQ2), bootstrapping was used with a stopping criterion that the average change is below 0.001. Regression analysis was performed using $R^{2}$ as goodness-of-fit metric to estimate the required number of participants to achieve similar results between conditions (included functions: normal, exponential, Weibull, logistic, Gompertz, gamma, beta, Rayleigh, lognormal and Pareto). For assessing the effects of personality traits (RQ3) and cognitive reflection (RQ4), comparison was performed between the upper and lower terciles of participants based on their scores on individual psychometric scales. Permutation tests with 20 permutations were performed to compare these effects to randomly selected baselines.
3.5.1 Structural analysis
While assessing the structural similarity of open card sorting results, we account for the uniqueness of categories created by individual participants by transforming the sorting data into similarity matrices. This is a standard approach for aggregate diverse categorizations in a consistent form that facilitates quantitative comparison [Macías and Culén, 2021, Sione Paea and Bulivou, 2024]. A similarity matrix is a grid that portrays how often cards were placed together pairwise. In the context where participants only see a subset of cards, similarity matrix values are relative to the number of participants to whom both cards were available simultaneously.
Let $S$ be $M\times M$ matrix, where $M$ is the number of cards and $N$ is the number of participants in the group. Each element $S_{i,j}$ represent the similarity score between cards $i$ and $j$. Function $p_{ij}(k)$ indicates whether participant $k$ paired cards $i$ and $j$ together (Equation 1) while function $s_{i,j}(k)$ indicates whether participant $k$ was shown both cards $i$ and $j$ (Equation 2).
$p_{ij}(k)=\begin{cases}1&\text{participant }k\text{ grouped cards }i\text{ and }j\text{ together}\\ 0&i=j\\ 0&\text{otherwise}\end{cases}$ (1)
[…]
Similarity matrix $S_{i,j}$ is then calculated as a ratio of participants who grouped cards $i$ and $j$ together to the number of participants who were shown both cards (Equation 3).
$S_{i,j}=\frac{\sum_{k=1}^{N}{p_{i,j}(k)}}{\sum_{k=1}^{N}{s_{i,j}(k)}}$ (3)
To enable quantitative comparison of similarity matrices, the Mantel test was used to calculate correlation between the matrices based on Spearman’s correlation [Guillot and Rousset, 2013]. As a complementary approach to assessing pairwise agreement within similarity matrices, we also used k-means to generate clusterings for a comparison of card organization groups and patterns suggested by the matrix. Optimal k was determined via the elbow method, individually for every resample in the case of the bootstrapping process. Adjusted Mutual Information (AMI) was applied as an evaluation technique, mitigating the potential bias tied to the Normalized Mutual Information (NMI) technique commonly used in literature [Mahmoudi and Jemielniak, 2024, Amelio and Pizzuti, 2017].
3.5.2 Linguistic analysis
Linguistic analysis of card sorting labels involved lemmatization of tokens and calculation of word frequencies. Chi-squared tests were used to assess similarity in the representations of words. To account for infrequent words that do not satisfy the test’s minimum assumption, Jaccard’s indices were also used as an aggregate assessment of the overlap of words in the generated labels. In addition to word distributions, we also analyzed the similarity of category labels based on their informativeness as a Gricean maxim—principle of effective language communication [Asada et al., 2022]. Considering the lesser informative value of common words, informativeness could be calculated based on inverted frequencies of words [Xiao et al., 2020]. Frequency analysis was supported by library wordfreq, based on a vast corpus of English text data [Speer, 2022].
3.5.3 Thematic analysis
Researchers qualitatively reviewed data to identify themes in categories representative of shared mental models. Standardizations were created based on the themes in category labels to account for categories with inconsistent labels (example from the dataset: phones, Telephony, Mobile & Tablets) [Paea et al., 2024]. To address researcher bias, two researchers with experience in conducting and analyzing card sorting studies collaboratively generated and consolidated standardizations, resolving the conflicts that emerged. Agreement scores of standardized categories $A_{c}$ were calculated as the card-wise sum of ratios of participants who sorted cards into included categories rather than a different category, divided by the number of included cards.
Let:
  • $m$ be the number of unique cards placed in the standardized category
  • $r$ be the number of respondents who contributed to the standardized category
  • $p_{k}$ be the number of respondents who placed card $k$ in the standardized category
Then, the agreeement score for the category $A_{c}$ is computed using Equation 4.
$A_{c}=\frac{100}{m*r}\sum_{k=1}^{m}p_{k}$ (4)
4 Results
Across both card sets, a higher number of cards elevated the time spent on card sorting. For the e-commerce card set, the difference was only moderate at median 633 seconds ($IQR:477\textendash 778$) in the Full-E group and 543 seconds ($IQR:404\textendash 783$) in the Subset-E group ($U(80)=953.5,z=1.48,p=.14$). The difference was more significant for the banking set at 748.5 seconds ($IQR:584.75\textendash 1324.25$) in the Full-B group and 539.5 seconds ($IQR:457.75\textendash 757$) in the Subset-B group ($U(80)=1126.5,z=3.14,p=.001,r=.35$).
A similar pattern was observed in the number of created categories. Within both groups Full-E and Subset-E recorded a consistent median at 5.5 ($IQR:4\textendash 7$) and no significant difference ($z=.43,p=.66$). However, between groups Full-B at 7.5 ($IQR:6\textendash 9$) and Subset-B at 6 ($IQR:4\textendash 7$), the number of created categories was significantly different ($U(80)=1107,z=2.95,p=.003,r=.33$).
The card set differences between groups also mediated relationships between subjective feedback measures. Among e-commerce groups Full-E and Subset-E, there was no significant difference in perceived spent time, label clarity, difficulty, focus, or even number of cards ($U(80)=815,z=0.14,p=.87$). However, with the more challenging banking card set, perception of the number of cards differed significantly ($U(80)=538.5,z=2.52,p=.005,r=.28$) between Full-B ($M=2.3$ out of 5, $SD=.73$) and Subset-B ($M=2.8$ out of 5, $SD=.56$). Perceived time spent also differed ($U(80)=628.5,z=1.65,p=.047,r=.18$) between Full-B ($M=2.8$ out of 5, $SD=.65$) and Subset-B ($M=3.1$ out of 5, $SD=.53$).
Personality traits and cognitive reflection test results had similar distribution between full and subset card sorting conditions, as shown in Figure 2. Therefore, effects of Extraversion ($U(160)=3661.5,p=.11$), Agreeableness ($U(160)=3337.5,p=.64$), Conscientiousness ($U(160)=3532,p=.25$), Negative Emotionality ($U(160)=3013,p=.52$), Open-Mindeness ($U(160)=3284,p=.77$) and Cognitive Reflection ($\chi^{2}(3,160)=5.83,p=.12$) can be compared between card set conditions (RQ3 and RQ4).
Figure 2: Overview of Big Five personality traits (a-e) and Cognitive Reflection (f) score distributions, which are similar between full card set and subset conditions.
4.1 Effect of randomized subsets (RQ1)
RQ1: How does card sorting with a randomized subset of cards affect its results?
Mantel tests confirm the alignment of results between full and subset card sorting conditions. Correlations of similarity matrices are strong for both card sets, e-commerce ($r=.93,p<.001$) and banking ($r=.78,p<.001$). Visual analysis of matrices shown in Figure 3 indicates this alignment as reasonably apparent. After applying k-means clusterings, obtained AMI scores reflect the association at 0.71 for e-commerce and 0.49 for banking.
Figure 3: Similarity matrices aggregating card sorting results for each experimental group, demonstrate strong parallels between full set (Full-E and Full-B) and subset (Subset-E and Subset-B) sorting conditions.
Word frequencies were analyzed in samples of 823 words for e-commerce and 820 for banking. The most frequent words were consistent across full and randomized subset card sorts. In e-commerce, they included “appliances”, “kitchen” and “electronic(s). In banking, the most frequent words were “insurance(s)”, “account(s)”, “saving(s)”. However, some words were represented at different frequencies, as shown in Table 3. Jaccard’s indices were 0.51 for e-commerce and 0.49 for banking. The informativeness of labels did not differ significantly between full and subset card sorts, neither in e-commerce $U(160)=921,z=1.16,p=.25$ or banking U(160)=766,z=0.33,p=.75U(160)=766,z=-0.33,p=.75.
Table 3: Words with different distributions between full set and subset conditions. Statistically significant differences are highlighted with *.
Card set Word Frequency (Full) Frequency (Subset) $\chi^{2}$
E-commerce gadget* 11 2 $\chi^{2}(1,823)=4.33,p=.037$
equipment* 11 2 $\chi^{2}(1,823)=4.33,p=.037$
small 0 8
large 0 8
Banking payment* 7 15 $\chi^{2}(1,820)=3.85,p=.049$
bank 2 8 $\chi^{2}(1,820)=3.63,p=.057$
investing 8 1 $\chi^{2}(1,820)=2.83,p=.092$
A summary of the thematic standardization of categories based card labels is shown in Table 4. Thematic standardizations had a similar level of low agreement $H(3,282)=3.57,p=.31$, indicating lack of internal consensus about cards belonging to these themes. For the e-commerce card set, using a randomized subset resulted in higher thematic variability, $\chi^{2}(1,472)=5.28,p=.022$. In the banking subset, which demonstrated innately higher variability even with the full card set, the difference in standardization count was not as significant $\chi^{2}(1,544)=2.90,p=.089$.
Table 4: Summary of standardization based on themes on category labels. Unstandardized categories could not be thematically grouped with other categories.
Condition Raw category count Standardization count Agreement (M) Unstandardized cat. count
Full-E 239 39 52.40% 10
Subset-E 233 59 53.81% 22
Full-B 304 93 47.37% 52
Subset-B 240 91 46.04% 49
Additionally, the most prevalent standardizations were represented more highly in full set card sorting, as seen in Figure 4 for e-commerce (e.g., Home/Household, Kitchen, Computers & Accessories) and Figure 5 for banking (e.g., Insurance, Accounts, Savings, Investments). Conversely, some of the less common standardizations were represented more frequently in the subset card sorting in both e-commerce (e.g., Personal, Gaming, Entertainment) and banking (e.g., Banking, Payments, Others).
Figure 4: Most frequent standardizations in the e-commerce conditions, based on number of participants who submitted original categories (threshold $\geq$ 4 participants).
Figure 5: Most frequent standardizations in the banking conditions, based on number of participants who submitted original categories (threshold $\geq$ 4 participants).
4.2 Number of participants (RQ2)
RQ2: How does the number of participants in card sorting with a randomized subset of cards affect its similarity to sorting with the full card set?
Results of bootstrapping within each experimental condition is shown Figure 6. The bootstrapped structural measures (correlation and AMI score) achieved stability in all conditions after 20 iterations. With the simpler e-commerce card set, when sorting the full list of cards, high average correlations were achieved even with small samples, as shown in Table 5. Sorting a card subset increased the number of participants needed to achieve similar correlation levels from 5 to 15 ($r\geq.90$) and from 8 to 22 ($r\geq.95$) respectively. The banking set had overall higher sample size requirements to achieve high correlations, but it was also further increased by the use of subsets, from 14 to 25 ($r\geq.90$) and from 22 to 31 ($r\geq.95$).
Figure 6: Within-condition bootstrapping of (a) correlation between similarity matrices and (b) AMI score between k-means clusterings. The number of iterations is 20.
Table 5: Determination sample sizes required to achieve strong correlation of similarity matrices or clustering similarity, based on within-condition bootstrapping. All correlations have $p<.001$.
Sample size requirement
Condition For $r\geq.90$ For $r\geq.95$ For $AMI\geq.7$ For $AMI\geq.8$
Full-E 5 ($r=.91$) 8 ($r=.95$) 4 ($AMI=.73$) 11 ($AMI=.80$)
Subset-E 15 ($r=.91$) 22 ($r=.95$) 14 ($AMI=.73$) 23 ($AMI=.81$)
Full-B 14 ($r=.90$) 22 ($r=.95$) 20 ($AMI=.73$) 35 ($AMI=.81$)
Subset-B 25 ($r=.90$) 31 ($r=.95$) 30 ($AMI=.71$) 37 ($AMI=.81$)
Following the assumption of Full-E and Full-B as the ground truth, we also performed between-condition bootstrapping as shown in Figure 7 to compare resampled Subset-E and Subset-B to the complete samples of Full-E and Full-B. In Subset-E, 11 participants are needed for correlation $r\geq.80$ and 24 for $r\geq.90$ ($r=.93$ between complete samples). In Subset-B, the requirement is 15 participants for $r\geq.60$ and 23 for $r\geq.70$ ($r=.78$ between complete samples). Regression models were applied to data points above 10 participants to accurately capture the more stable trends at increasing values. Weibull cumulative distribution was the best fit at very high $R^{2}=0.993$. This model estimates that to achieve $r\geq.90$, 82 participants would be needed. AMI scores demonstrate similar patterns.
Figure 7: Between-condition bootstrapping of (a) correlation between similarity matrices and (b) AMI score between k-means clusterings. The number of iterations is 20.
4.3 Personality traits (RQ3)
RQ3: How do participants’ personality traits affect their card sorting results?
The strength of correlation between similarity matrices obtained from high and low personality trait subgroups varied depending on card sorting conditions. The strength of the correlations followed a consistent pattern across all Big Five traits—very strong in Full-E, moderate-to-strong in Subset-E and Full-B, and weak in Subset-B (see Table 6). These effects are comparable to the random baseline as evaluated through permutation tests.
Table 6: Comparison of correlations between high and low scorers in Big Five personality traits, Cognitive Reflection, and randomly selected permutations.
Condition Big Five traits Cognitive Reflection Chance (20 permutations)
Full-E 0.880.930.88-0.93 0.91 $M=0.91,SD=0.02$
Subset-E 0.660.720.66-0.72 0.55 $M=0.69,SD=0.03$
Full-B 0.670.710.67-0.71 0.67 $M=0.68,SD=0.03$
Subset-B 0.320.400.32-0.40 0.24 $M=0.35,SD=0.03$
In a small number of group-specific instances, differences between high and low personality trait scores corresponded with significant differences in other observed measures (see Table 7). Participants with low Conscientiousness found it harder to focus in Subset-B and Subset-E than those with high Conscientiousness. In condition Full-B, more Open-Minded participants were more comfortable with the higher Card Count and created card labels with higher Informativeness. In Subset-B, higher Negative Emotionality was also associated by more Informative labels.
Table 7: Significant differences between participants with high and low Big Five traits, N=26.
Condition Trait Measure Low trait group High trait group U-test
Subset-E Conscientiousness Perceived Focus $M=4.54$
$SD=0.66$
$M=5$
$SD=0$
U=52,z=1.67U=52,z=-1.67
$p=.017,r=.33$
Subset-B Conscientiousness Perceived Focus $M=4$
$SD=1.15$
$M=4.85$
$SD=0.38$
U=48,z=1.87U=48,z=-1.87
$p=.029,r=.37$
Subset-B Negative Emotionality Informativeness $M=18.67$
$SD=3.89$
$M=26.28$
$SD=6.92$
U=28,z=2.90U=28,z=-2.90
$p=.004,r=.57$
Full-B Open-Mindedness Perceived Card Count $M=1.92$
$SD=0.76$
$M=2.62$
$SD=0.65$
U=42.5,z=2.15U=42.5,z=-2.15
$p=.021,r=.42$
Full-B Open-Mindedness Informativeness $M=18.67$
$SD=3.35$
$M=23.56$
$SD=5.9$
U=37,z=2.44U=37,z=-2.44
$p=.016,r=.48$
Within low and high personality trait scorer groups, specific distinctions could be found in measures that are significantly different between the full card set and subset. As shown in Table 10, personality-linked differences between the full card set and subset were more pronounced in the Banking card set, with several attributed to each Big Five personality trait. This demonstrates several types of potential influence of personality on card sorting behavior and corresponding user attitudes as interpreted in 5.1 Implications.
To explore specific qualitative properties of categories created by participants with different personalities as the basis for future hypotheses, we conducted subgroup analysis of themes. Although the lack of statistical significance in subgroups prevents differences in categories (shown through standardization counts in Table 11) from being confidently attributed to personality, it indicates potential patterns and inconsistencies between groups (see interpretation in 5.1 Implications).
4.4 Cognitive reflection (RQ4)
RQ4: How does participants’ cognitive reflection affect their card sorting results?
The strength of correlations between similarity matrices from high and low CRT scorers was dependent on the experimental condition. It followed a similar pattern of personality subgroups, albeit weaker for subset groups: very strong in Full-E, moderate in Subset-E, moderate-strong in Full-B and weak in Subset-B (see Table 6).
Significant differences between Full-B and Subset-B that were not present among low-scorers were identified among high CRT scorers. Affected variables include Time Spent ($M_{Full}=1013.23,M_{Subset}=555.85,U(26)=139,z=2.79,p=.006$), Category Count ($M_{Full}=8.31$, $M_{Subset}=4.62$, $U(26)=152,z=3.46,p<.001$) and Perceived Time Spent ($M_{Full}=2.62$, $M_{Subset}=3.15$, U(26)=49.5,z=1.79,p=.016U(26)=49.5,z=-1.79,p=.016). Based on thematic analysis of subgroups shown in Table 11, high CRT scorers also created fewer thematic standardizations.
5 Discussion
Findings from our experiment highlight the substantial impact of the investigated design factors and sample variables, with implications for result validity and potential for more robust methodologies for planning, conducting and interpreting data from card sorting in UX research.
5.1 Implications
RQ1: How does card sorting with a randomized subset of cards affect its results?
As long as card sorting analysis focuses primarily on aggregated results in the form of similarity matrices or clusterings, our findings validate the use of randomized subsets as more efficient substitutes for the full set of cards. High correlations and AMI scores demonstrate that mental models constructed from card pairings can be highly similar. However, the simpler e-commerce card set involving physical objects yielded higher similarity compared to the complex and less familiar banking card set. This could be seen as evidence that randomly selecting from unfamiliar contents at the context could introduce higher variability in represented mental models due to inconsistent perspectives.
Linguistic and thematic analysis—through differences in word and theme frequencies—support the presence of effects that become more pronounced during qualitative analysis of specific created categories. While Jaccard’s indices correspond with natural variability in category name data, with descending frequencies in agreement with Zipf’s law [Aitchison et al., 2016] regardless of whether random subsets are used, the exact concepts represented by created categories can become altered (e.g. focus on size of products instead of functional themes such as kitchen gadgets, music gadgets, cleaning gadgets). As shown in both Figure 4 and Figure 5, the most prevalent mental models can appear as less universally adopted while more niche models manifest as more popular, creating a bias of falsely amplified dissent. Therefore, analysis of specific categories from card sorting with randomized subsets can contain noise and be less representative of the real prevalence rate of specific mental models among users (e.g., as a measure of intuitiveness to predict their reliability as labels in information architectures).
If the intent behind a card sorting activity is to examine mental models from more varied perspectives instead of just focusing on the most intuitive ones, random subsets can function as a catalyst for increased variability. This method could approximate variable data from internally diverse target groups in practical scenarios where recruiting larger or more varied samples is not feasible. As evidenced by increased standardization and unstandardized category count in the e-commerce card set and insignificant change in the banking set (see Table 4), card sets that are familiar and easy to sort intuitively may benefit more from this approach.
RQ2: How does the number of participants in card sorting with a randomized subset of cards affect its similarity to sorting with the full card set?
Our analysis corroborates the findings by Lantz et al. [2019] that in standard card sorting with a full card set, 10-15 can be the optimal number of participants [Lantz et al., 2019]. However, they also demonstrate that a higher number of participants is needed when sorting a randomized subset. Furthermore, card set difficulty and unfamiliarity can increase the requirement to achieve balanced correlation. Considering subsets that constitute 60% of the original set, 25 participants ($r=.90$) and 35 participants ($r=.95$) could be considered as optimal for card sorting of reasonable difficulty. For thematically simpler card sorts, such as those involving everyday physical objects, 15-25 participants can also be sufficient.
To offer a guideline for calculating the number of participants in studies that utilize randomized subsets of different sizes, we present a simplified mathematical model as a function that fits optimal sample sizes at 60% and 100% subset thresholds. Reasonable expectations suggest that as the size of subsets seen by participants approaches zero, patterns will deviate more due to differences in perceived context, thus making larger samples necessary. Therefore, the sample size determination formula for card sorting with randomized card subsets is proposed as a logarithmic decay function shown in Equation 5, where $|C|$ is the size of the full set of all cards and $|S\subseteq C|$ is the size of randomized subsets.
$N=15-90\log{\frac{|S\subseteq C|}{|C|}}$ (5)
The minimum threshold could be significantly higher (80-90) if the goal was adjusted beyond just internal result stability in order to achieve results that highly correlate with standard card sorting without randomized subsets. Due to differences pertaining to categories created through card sorting with and without randomized subsets however, it may be more appropriate to acknowledge standard and randomized subset card sorting as distinct approaches with individual pros and cons. Higher variability in categories can be expected with randomized subsets, whether or not a larger sample achieves more consistent pairwise correlation.
RQ3: How do participants’ personality traits affect their card sorting results?
Personality can affect how people approach card sorting. Big Five personality traits were found to be mostly aligned with behavioral and emotional characteristics typically associated with corresponding traits. This effect was more pronounced in the more challenging banking card set (Full-B and Subset-B).
Negative Emotionality. Vulnerability to negative emotions was reflected as higher sensitivity toward the amount of time spent during the activity. In the challenging Banking card set, high Negative Emotionality drove participants to create more categories with more unique labels when sorting the full set, potentially due to the stress of the cognitively complex exercise. Qualitative analysis of categories by this group revealed more generic low-quality categories, as well as smaller categories linked through obvious associations, which could have resulted from the synergy between anxiety (to complete large card sorting in a timely manner) and low-level sorting strategies becoming more plausible due to the higher number of cards (reduced need to mentally process more distant logical links). Low Negative Emotionality associated with significantly more time spent by engaging with the full card set might indicate a lower tendency to satisfice when encountering a higher number of cards.
Extraversion. Unlike introverts, extroverts spent significantly more time on Full-B than on Subset-B, which could be attributed to their sociability and high energy driving them to dedicate more attention to building their mental model of the cards. Previous studies that associated extroversion with better flow, even in solitary activities [Tse et al., 2025] might corroborate this. Meanwhile, introverts were more inclined to rate the 50-card activity as taking long and view the card count as closer to ”too many”.
Agreeableness. As a trait associated with prosocial behavior [Habashi et al., 2016], it is intuitive that when participants with high Agreeableness were given more cards, they created significantly more categories in an effort to be helpful. Meanwhile, disagreeable participants reported the full set activity as significantly more difficult. This could be due to various factors, such as lack of internal motivation, or less prominent tendency toward telling prosocial lies [Reinhardt et al., 2024].
Conscientiousness. Participants with low conscientiousness found it significantly more difficult to focus, which can be linked to the nature of card sorting as a complex task. Full-B taking them significantly more time to complete and resulting in significantly more categories than Subset-B can be attributed to the higher difficulty in identifying logical patterns. Participants with high conscientiousness also found it more difficult to focus with a full set of cards, which could be connected to their higher perceptiveness toward the mental challenge, since they are in the habit of solving tasks in a systematic and orderly manner.
Open-Mindedness. Highly Open-Minded participants—characterized by being imaginative and curious—demonstrated this trait by creating more informative category labels. Open-Mindedness was also linked to attitudes to the number of cards in the challenging Full-B condition—low scorers viewed it negatively while high-scorers were neutral. Low-scorers completed Subset-E significantly faster than Full-E. High-scorers spent significantly more time on Full-B than on Subset-B. These findings could indicate that low-scorers are less likely to spend time cognitively processing simpler card sorting tasks and are faster to proceed with sorting. High-scorers might tend to expend significant effort on tasks that provide deeper opportunities for creative self-expression.
Consistency in correlation patterns between high and low scorers across all card sorting conditions implies that all personality traits had similar impact on obtained card pairings. Given the similarity to correlations obtained when permuting participants randomly, we can assert that the effect of personality was minimal, with lower correlations caused by variability of data in respective card sorting methods rather than personality. With a sufficient number of participants and realistic representation of varied personality trait combinations, similar patterns can be expected to emerge. Furthermore, even recruitment methods skewed in favor of specific personalities (e.g., introverts) may also result in comparable similarity matrices.
However, personality can affect user engagement (time spent, category count) and subjective experience during card sorting. Furthermore, explorative analysis might signal possible impacts of personality on the quality and thematic properties of created categories (see Table 11). Interpreted, this could potentially include phenomena such as participants high in:
  • Agreeableness generating more themes within challenging larger card sets (Full-B) because of their helpfulness,
  • Conscientiousness being more meticulous, systematically creating categories that are the best match the set of cards given to them, thus creating more themes across varied complex subsets (Subset-B),
  • Negative Emotionality reacting to potentially stressors to generate more unique themes while sorting subsets (Subset-B), while low scores might contribute to the discovery of more themes within a larger set of cards (Full-B) because of better coping ability,
  • Open-Mindedness expressing more creative themes in complex card sets (Full-B and Subset-B), yet being less influenced by reduced context is simpler card sets (Subset-E), generating a similar number of themes as they would if given the full context (Full-E).
The concept of targeting specific personalities to capitalize on personality traits (e.g., high agreeableness for more careful categorization in complex card sorting, open-mindedness for more imaginative categories) or mitigate issues (e.g., rushing in complex studies due to high negative emotionality) can be seen as promising, although it could also raise concerns about bias (social desirability, categories that are designer-like rather than intuitive). However, since overarching patterns in similarity matrices should be consistent across personalities, with only themes, effort, and sorting experience being affected significantly, an argument could be made for screening based on desirable personalities to obtain categories that match a personality profile.
RQ4: How does participants’ cognitive reflection affect their card sorting results?
In contrast to personality traits, CRT subgroups correlated in Subset-E and Subset-B more weakly than correlations between subgroups obtained through permutation testing. Therefore, cognitive reflection likely plays a role in the cognitive processes tied to randomized subset card sorting. A possible explanation is that higher cognitive reflection allowed participants to systematically assess the context. Either through better intuition [Thompson et al., 2018] or reflective reasoning, the cognitively-reflective disposition may enable participants to mentally fill in the logical gaps left by the random selection of card subsets. This interpretation would explain why high scorers can generate fewer (more consistent) themes in a complex card set (Subset-B), with fewer categories created to match the specific randomizations.
Therefore, when conducting card sorting with randomized subsets, focusing on participants with high cognitive reflection during recruitment or data analysis may contribute to higher quality of findings. CRT or its alternatives such as CRT-2 and CRT-V [Thomson and Oppenheimer, 2016, Sirota et al., 2021] are conveniently short. As such, they may be practical to efficiently diagnose cognitive reflection deficiency. To obtain participants who spend more effort on sorting cards (time, category count), administering CRT may also be beneficial for standard card sorting.
5.2 Limitations and threats to validity
The presented findings could be limited by factors that stem from the applied experimental and analytical methods. Internal validity threats, which are standard for between-subjects design (potential differences between samples) were addressed through stratified random sampling aimed at mitigating systematic effects of confounding variables. External validity could be limited geographically and culturally due to the recruitment of UK residents. The data is linked to two card sorting studies; therefore, the findings could be limited to contextual characteristics specific to these studies and effects that they may exert on users. The online and remote administration of the experiment ensures its high ecological validity tied to naturalistic settings.
Construct validity was maintained through the use of established psychometric tools (BFI-2-XS, CRT) and standard card sorting measures and analytical tools. Conclusion validity was achieved through sample size 40 participants per group, which is higher than typically suggested number of participants to obtain reliable results in card sorting [Pechlevanoudis et al., 2023] and determined as sufficient for medium effect size ($r=0.3$) via a priori test (see 3.1 Participants). Results of subgroup analysis (in RQ3 and RQ4) should be understood as explorative, with focus on large effects given the smaller subgroup size. To increase the credibility of conclusions through dialogue and mitigate researcher bias, findings were analyzed collaboratively by two researchers and a third resolved disagreements.
The proposed sample size determination formula for card sorting with randomized subsets extrapolates upon findings from four experimental conditions where either 60% or 100% of all cards were presented to participants. As such, it may not accurately fit other ratios. Empirical examination using different subset size ratios may reveal functions other than logarithmic decay to be a better fit. For card subsets of small size (relative to the full card set), even results collected from large samples may become too noisy or they may diverge from standard card sorting more significantly, since there is a higher chance of key cards not being included.
5.3 Future work
To increase the generalizability of findings regarding the effects of randomized subsets in real-world card sorting studies, future research could aim for replicability in the context of different domains, using different full-card-set-to-subset ratios, involving different user groups, levels of familiarity and complexity of the cards as well as their general contents.
The effects of personality and cognitive reflection on card sorting could be further assessed to enable validation of the proposed explanations based on the findings of subgroup analysis. For example, a study involving larger samples representing low and high scorers in negative emotionality could examine thematic patterns in more detail to determine how stressors in card sorting design may contribute to the quality of generated categories (e.g., generic vs. idiosyncratic, goal-oriented vs. model-based, surface-level vs. deep-level). Self-reported information about satisficing, socially desirable responding or anxiety could be collected to increase the credibility of explanations.
Studying the relationship between navigation in hierarchical information structures and Big Five traits or Cognitive Reflection might offer further insights about the links between mental models, psychometric constructs and navigation behavior. Selecting participants with preferable traits for card sorting because of their higher emotional stability, or a detail-oriented or creative approach may appear appealing to obtain high-quality data. However, the effects on navigation performance of other personality groups in resulting information architectures should be studied to prevent designing information architectures that might be less accessible to specific personality types.
Since card sorting is a time-demanding generative exercise, further increasing its length by preceding it with a personality assessment questionnaire to increase its accuracy might be seen as impractical even with shortened adaptations like the BFI-2-XS. Future research could apply machine learning with the aim to predict the participant’s personality based on behavioral cues as a tool for AI-assisted analysis.
It was a precondition for this study to ensure the validity and feasibility of multifaceted analysis of card sorting results. As such, our experiment was designed to minimize attrition through strategies like participant compensations and a reasonable number of cards in all conditions. However, future experiments investigating the effects of personality could focus on attrition. For illustration, persons with high negative emotionality or extraversion could be more inclined to abandon challenging card sorting studies, similarly to a risk demonstrated by Hansson et al. [2018] in longitudinal studies.
Future studies could examine which cards (by order or label properties) are placed into categories first, or which ones result in the creation of new categories. Sorting strategies, such as sorting cards in the same order as cards are presented, or prioritizing cards by perceived relationships, could be assessed. If different strategies are preferred by different personalities or with subset randomization, they could be a factor for differences between resulting categories. The point of saturation (when sorting more cards results in no further creation, renaming or removal of categories) could be explored for deeper understanding. Potential applications could include prediction of the appropriate number of presented cards and participants in different contexts, or the design of a hybrid method where a portion of cards could be fixed because of their significance, while the other portion could be randomly rotating.
To assess the potential for variability in randomized subset card sorting to realistically simulate variability in internally diverse populations, studies could compare it to standard card sorting with large samples that capture high diversity through strategies like stratified sampling to cover a variety of user types and demographics.
6 Conclusion
In this article, we examined the validity of the common practice in card sorting: using randomized card subsets to organize a high number of items while maintaining reasonable procedure duration. Previously, the field lacked research that would comprehensively compare this approach with results yielded by standard open card sorting with all cards. For robust understanding, we conducted an experiment with 160 participants across 2 card sorting studies of different complexity, including personality inventory and cognitive reflection assessment instruments for further insights. We analyzed data from multiple perspectives, including their results, optimal number of participants as well as subgroup exploration to assess the potential impacts of Big Five personality traits and Cognitive Reflection Test scores between conditions.
We demonstrated that user researchers can rely on card sorting with randomized subsets that comprise 60% of the complete set to obtain similar pairwise similarity matrices to the standard approach. However, thematic and linguistic analyses indicate significant differences in concrete categories. As a result of higher variability, randomized subset card sorting was discovered to require more participants (25-35 at 60% of the full card set) to achieve stable results. To address this, we proposed a sample size determination formula in which the total number of cards and the size of the subset can be entered as inputs. As a common factor between personality and cognitive reflection, psychological factors they can affect card sorting in a variety of manners, including participant behavior, effort, as well as structural and thematic properties of results. We view these relationships as a salient topic for further study aimed at improving methodological rigor and efficiency of card sorting, as well as the quality of its results for better understanding mental models in practical scenarios.
Acknowledgements
This work was supported by the Slovak Research and Development Agency under Contract No. APVV-23-0408, and co-financed by the Cultural and Educational Grant Agency of the Slovak Republic (KEGA) under Grant No. KG 014STU-4/2024. We would like to thank UXtweak j.s.a. for their generous financial contribution to this research, as well as the UXtweak Research team for their technical and expert support.
Data availability statement
Supplementary data and materials, including data files containing participants’ results, analysis scripts and supplementary experiment materials are available in the paper repository at https://github.com/cardsort-research/card-sorting-design
Author contributions
Eduard Kuric: Writing – review & editing, Writing – original draft, Validation, Supervision, Resources, Project administration, Methodology, Investigation, Funding acquisition, Formal analysis, Conceptualization. Peter Demcak: Writing – review & editing, Writing – original draft, Visualization, Validation, Methodology, Investigation, Formal analysis. Matus Krajcovic: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation.
Disclosure statement
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix A Card sets
Card sets used in e-commerce (see Table 8) and banking (Table 9) variants.
Table 8: List of cards used in the e-commerce variant, containing electrical devices and appliances.
E-commerce variant card set
Air Conditioners Gaming Consoles Printers
Air Purifiers Graphic Cards Processors
Cameras Hair Dryers Radiators
Chargers Headphones Radios
Coffee Makers Hobs Scanners
Computer Mice Irons Shavers
Cookers Juicers Smart Watches
Deep Fryers Keyboards Smartphones
Desktop Computers Kitchen Scales Speakers
Dishwashers Laptops Stand Mixers
Drones Microphones Tablets
E-Readers Microwave Ovens Televisions
Electric Kettles Monitors Trimmers
Electric Toothbrushes Musical Instruments Vacuum Cleaners
Fans Ovens VR Headsets
Freezers PC Games Washing Machines
Fridges Phone Cases
Table 9: List of cards used in the banking variant, containing more abstract banking-related concepts.
Banking variant card set
Account Statements Financial Education Webinars Online Purchases
Balance Transfer Fixed Deposit Other Borrowing Options
Bereavement Support Fixed Term Savings Overdraft Protection
Bonds Fixed-rate Credit Card Personal Loan Calculator
Budgeting Tools and Resources Fraud Alerts and Monitoring Personal Pension
Car Insurance Renewal Fund Performance Personalised Financial Advice
Cardless Cash Withdrawal Google/Apple Pay Remortgage Options
Cheque Payments Insurance Coverage Calculator Retirement Planning Calculator
Claims Interest Rates Rewards Programme
Credit Score International Payments Savings Options
Current Accounts International Trading Self-Invested Personal Pension (SIPP)
Daily Savings Investment Accounts Stock Trading and Shares
Data and Privacy Control ISA Accounts Student account
Debit Cards Life Insurance Coverage Student Loan Repayment
Debt Management Services Mobile Banking App Travel Insurance
Digital Confidence Mortgage Calculator Trustee Banking
Financial Assistance Mortgage Rates
Appendix B Supplementary sub-group analysis tables
Significant differences between full and subset card sets (Table 10) and subgroup theme analyses for high standardization differences (Table 11).
Table 10: Significant differences between full card set and subset within groups of high and low Big Five trait scorers, N=26.
Card set Trait Score Measure Full variant Subset variant U-test
E-commerce Conscientiousness High Perceived Focus $M=4.62$
$SD=5$
$M=0.65$
$SD=0$
U=58.5,z=1.33U=58.5,z=-1.33
$p=.037,r=.26$
E-commerce Negative Emotionality High Perceived Time Spent $M=3.15$
$SD=2.69$
$M=0.38$
$SD=0.63$
$U=118.5,z=1.74$
$p=.035,r=.34$
E-commerce Open-Mindedness Low Time Spent $M=711.15$
$SD=535.85$
$M=197.44$
$SD=237.50$
$U=126,z=2.13$
$p=.035,r=.42$
Banking Extraversion High Time Spent $M=972.46$
$SD=609.69$
$M=491.77$
$SD=226.80$
$U=127,z=2.18$
$p=.031,r=.43$
Banking Extraversion Low Perceived Time Spent $M=2.62$
$SD=3.31$
$M=0.65$
$SD=0.48$
U=38,z=2.38U=38,z=-2.38
$p=0.008,r=.47$
Banking Extraversion Low Perceived Card Count $M=2.15$
$SD=2.92$
$M=0.80$
$SD=0.64$
U=42.5,z=2.15U=42.5,z=-2.15
$p=.021,r=.42$
Banking Agreeableness High Category Count $M=8.38$
$SD=6.38$
$M=2.33$
$SD=1.98$
$U=127,z=2.18$
$p=.029,r=.43$
Banking Agreeableness Low Perceived Difficulty $M=3.08$
$SD=3.69$
$M=0.76$
$SD=0.85$
U=48,z=1.87U=48,z=-1.87
$p=.045,r=.37$
Banking Conscientiousness Low Time Spent $M=854$
$SD=561.15$
$M=347.21$
$SD=132.62$
$U=138,z=2.74$
$p=.007,r=.54$
Banking Conscientiousness Low Category Count $M=8.15$
$SD=5.38$
$M=2.27$
$SD=2.06$
$U=140,z=2.85$
$p=.004,r=.56$
Banking Conscientiousness High Perceived Focus $M=3.85$
$SD=4.85$
$M=0.99$
$SD=0.38$
U=34,z=2.59U=34,z=-2.59
$p=0.004,r=.51$
Banking Negative Emotionality Low Time Spent $M=1147.08$
$SD=633.31$
$M=613.89$
$SD=228.25$
$U=124,z=2.03$
$p=.046,r=.40$
Banking Negative Emotionality High Category Count $M=7.46$
$SD=5.62$
$M=2.47$
$SD=2.14$
$U=125.5,z=2.10$
$p=.035,r=.41$
Banking Negative Emotionality High Informativeness $M=19.06$
$SD=26.28$
$M=7.09$
$SD=6.92$
U=34,z=2.59U=34,z=-2.59
$p=.010,r=.51$
Banking Open-Mindedness High Time Spent $M=1019.23$
$SD=578.38$
$M=492.07$
$SD=243.31$
$U=139,z=2.79$
$p=.005,r=.55$
Banking Open-Mindedness Low Perceived Card Count $M=1.9$
$SD=2.92$
$M=0.76$
$SD=0.49$
U=27,z=2.95U=27,z=-2.95
$p=.001,r=.58$
Table 11: Subgroup analysis of themes (card standardizations and unstandardized categories) for conditions with highest differences in standardization count (>6>6).
Trait Variant Group Standardization count Agreement (M) Raw category count Unstandardized cat. count
Agreeableness Full-B Low 38 34.26 95 9
High 51 38.09 109 24
Conscientiousness Subset-B Low 38 33.76 70 10
High 50 34.01 84 20
Negative Emotionality Full-B Low 48 35.06 106 24
High 40 36.16 97 10
Subset-B Low 44 33.7 85 11
High 51 33.69 73 23
Open-Mindedness Subset-E Low 37 41.38 78 8
High 27 37.11 70 2
Full-B Low 45 35.04 103 14
High 55 38.16 101 26
Subset-B Low 51 33.8 85 19
High 42 35.01 60 16
Cognitive Reflection Subset-B Low 51 33.8 85 19
High 42 35.01 60 16

### v2

@@
-Citation: Kuric E., Demcak P., Krajcovic M. (2025) Talking Surveys: How Photorealistic Embodied Conversational Agents Shape Response Quality, Engagement, and Satisfaction
+Card Sorting with Fewer Cards and the Same Mental Models? A Re-examination of an Established Practice
 Abstract
@@
 Representation of specific profiles may also be beneficial, leading us to investigate the effects of personality and cognitive reflection.
-This is a topic yet to be examined for card sorting in general, in spite of personality and cognitive reflection having established effects on user behavior, experience and perceived usability [Hattie Wilczewski and Kortum, 2024, Liapis et al., 2019, Özbek et al., 2014, Mosleh et al., 2021], as well as relationships with mental models in other contexts [Vˆırgă et al., 2014, Fisher et al., 2012].
+This is a topic yet to be examined for card sorting in general, in spite of personality and cognitive reflection having established effects on user behavior, experience and perceived usability [Hattie Wilczewski and Kortum, 2024, Liapis et al., 2019, Özbek et al., 2014, Mosleh et al., 2021], as well as relationships with mental models in other contexts [Vîrgă et al., 2014, Fisher et al., 2012].
 Therefore, key contributions of this research consist of:
@@
 Even though alternatives exist, such as the six-factor HEXACO—which offers higher orthogonality along with potential theoretical and practical advantages [Ashton and Lee, 2007, Thielmann et al., 2022]—the Big Five remains the most prolific personality model at use today.
-Big Five consists of five traits: Negative Emotionality, Extraversion, Agreeableness, Conscientiousness and Open-Mindedness, associated with the following characteristics [Zekry and McKee, 2023, Tatnall, 2020]:
+Big Five consists of five traits: Negative Emotionality, Extraversion, Agreeableness, Conscientiousness and Open-Mindedness, associated with the following characteristics [D.
+A.
+Zekry and G.
+T.
+McKee (2023), 7]:
   • Negative Emotionality. Susceptibility or resistance to negative feelings like fear or stress.
@@
 The pre-study questionnaire contained a demographic questionnaire to monitor the sample’s distribution and a Cognitive Reflection Test (CRT) to measure cognitive capacity.
-The post-study questionnaire contained the personality assessment questionnaire (BFI-2-XS) [Soto and John, 2017a].
+The post-study questionnaire contained the personality assessment questionnaire (BFI-2-XS) [Soto and John, 2017b].
 Both questionnaires also incorporated an attention check question to verify participants’ engagement and effort during the study.
@@
 3.3.3 Big Five Inventory-2
-To minimize potential effects of questionnaires on engagement and thus on user behavior during card sorting, we used the extra short form of the Big Five Inventory-2 (BFI-2-XS) [Soto and John, 2017a] to assess Big Five personality traits.
-While the standard BFI-2 is a robust instrument for measuring the five personality traits and their respective three facets [Soto and John, 2017b], it contains 60 Likert scale items, which is considerable when the efficiency of data collection is a concern.
+To minimize potential effects of questionnaires on engagement and thus on user behavior during card sorting, we used the extra short form of the Big Five Inventory-2 (BFI-2-XS) [Soto and John, 2017b] to assess Big Five personality traits.
+While the standard BFI-2 is a robust instrument for measuring the five personality traits and their respective three facets [Soto and John, 2017a], it contains 60 Likert scale items, which is considerable when the efficiency of data collection is a concern.
 Card sorting itself is already an effort-intensive process.
 As such, the 15-item BFI-2-XS was chosen to prevent fatigue from participants.
-In spite of its short length, BFI-2-XS captures the Big Five traits—Extraversion, Conscientiousness, Open-Mindedness, Agreeableness and Negative Emotionality—with 80% of the variability of BFI-2 [Soto and John, 2017a].
+In spite of its short length, BFI-2-XS captures the Big Five traits—Extraversion, Conscientiousness, Open-Mindedness, Agreeableness and Negative Emotionality—with 80% of the variability of BFI-2 [Soto and John, 2017b].
 3.3.4 Cognitive Reflection Test
@@
 Across both card sets, a higher number of cards elevated the time spent on card sorting.
-For the e-commerce card set, the difference was only moderate at median 633 seconds ($IQR:477\textendash 778$) in the Full-E group and 543 seconds ($IQR:404\textendash 783$) in the Subset-E group ($U(80)=953.5,z=1.48,p=.14$).
-The difference was more significant for the banking set at 748.5 seconds ($IQR:584.75\textendash 1324.25$) in the Full-B group and 539.5 seconds ($IQR:457.75\textendash 757$) in the Subset-B group ($U(80)=1126.5,z=3.14,p=.001,r=.35$).
+For the e-commerce card set, the difference was only moderate at median 633 seconds (IQR:477778IQR:477–778) in the Full-E group and 543 seconds (IQR:404783IQR:404–783) in the Subset-E group ($U(80)=953.5,z=1.48,p=.14$).
+The difference was more significant for the banking set at 748.5 seconds (IQR:584.751324.25IQR:584.75–1324.25) in the Full-B group and 539.5 seconds (IQR:457.75757IQR:457.75–757) in the Subset-B group ($U(80)=1126.5,z=3.14,p=.001,r=.35$).
 A similar pattern was observed in the number of created categories.
-Within both groups Full-E and Subset-E recorded a consistent median at 5.5 ($IQR:4\textendash 7$) and no significant difference ($z=.43,p=.66$).
-However, between groups Full-B at 7.5 ($IQR:6\textendash 9$) and Subset-B at 6 ($IQR:4\textendash 7$), the number of created categories was significantly different ($U(80)=1107,z=2.95,p=.003,r=.33$).
+Within both groups Full-E and Subset-E recorded a consistent median at 5.5 (IQR:47IQR:4–7) and no significant difference ($z=.43,p=.66$).
+However, between groups Full-B at 7.5 (IQR:69IQR:6–9) and Subset-B at 6 (IQR:47IQR:4–7), the number of created categories was significantly different ($U(80)=1107,z=2.95,p=.003,r=.33$).
 The card set differences between groups also mediated relationships between subjective feedback measures.